solve the quadrtic equations by completeing the square, the square root property or quadratic formula.
a. (x+3)^2-81=0
b. 3x^2+5x=2
c. (x+1)(x-2)=18
d. 2x^2-4x+5=0

Answer by sachi(548)

a. (x+3)^2-81=0
or (x+3)^2=81
or x+3=sqrt 81=+9 or -9
or x=9-3 or -9-3=6 or -12
b. 3x^2+5x=2
or 3x^2+5x-2=0
or 3x^2+6x-x-2=0
or 3x(x+2)-1(x+2)=0
or (x+2)(3x-1)=0
or x=-2 or 1/3
c. (x+1)(x-2)=18
or x^2-2x+x-2-18=0
or x^2-x-20=0
or x^2-5x+4x-20=0
or x(x-5)+4(x-5)=0
or (x-5)(x+4)=0
or x=5 or -4
d. 2x^2-4x+5=0 solution by completing square
or 4*2*(2x^2-4x+5)=0 multiplying 4a on both sides
or 16x^2-32x+40=0
or (4x)^2-2*4x*4+4^2-4^2+40=0 adding & subtracting 4^2
or (4x-4)^2=16-40=-24
or 4x-4=+/-sqrt(-24)=+/-2sqrt(-6)=+/-2isqrt6 where i=sqrt(-1)
or 4x=4+/-2isqrt6
or x=1+/-0.5i sqrt6